### v1

Abstract.
The increasing affordability of robot hardware is accelerating the integration of robots into everyday activities. However, training a robot to automate a task typically requires physical robots and expensive demonstration data from trained human annotators. Consequently, only those with access to physical robots produce demonstrations to train robots. To mitigate this issue, we introduce EVE, an iOS app that enables everyday users to train robots using intuitive augmented reality visualizations without needing a physical robot. With EVE, users can collect demonstrations by specifying waypoints with their hands, visually inspecting the environment for obstacles, modifying existing waypoints, and verifying collected trajectories. In a user study ($N=14$, $D=30$) consisting of three common tabletop tasks, EVE outperformed three state-of-the-art interfaces in success rate and was comparable to kinesthetic teaching—physically moving a real robot—in completion time, usability, motion intent communication, enjoyment, and preference ($mean_{p}=0.30$). We conclude by enumerating limitations and design considerations for future AR-based demonstration collection systems for robotics.
augmented reality, demonstration collection, imitation learning
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: ; April, 2024; ccs: Human-centered computing Mixed / augmented realityccs: Computer systems organization Robotics
1. Introduction
Figure 1. EVE enables everyday users to collect data to train a real robot using intuitive augmented reality (AR) visualizations without needing a physical robot. Using our interface, the user can move the AR robot by setting waypoints with the hand, visually inspect the real world to avoid obstacles, modify the collected trajectory by reverting to a previous waypoint, and verify the collected trajectory by replaying the task with the AR robot. See supplementary materials for the videos of real-world deployment.
The decreasing costs of robot components and systems are accelerating the integration of robots into everyday life. With companies offering robotic arms at a few hundred dollars, portable, lightweight, and mobile robots are being trained to assist individuals with a wide range of everyday tasks (Elephant Robotics, 2023; Robot, 2023; Shaw et al., 2023a). However, training a robot to automate a task often requires expensive manually curated human demonstrations. Annotators are hired and trained to kinesthetically move a robot’s joints and gripper to a sequence of configurations to accomplish a task. For example, to train a robot to clean a table, annotators move the gripper towards a cloth, close the gripper to grasp the cloth, and then begin moving it around to wipe the table.
As such, there are no mechanisms for everyday users, who are considering purchasing a robot, to train robots without first purchasing an expensive robot and learning to move bulky machinery to produce effective demonstrations. Consequently, only a limited number of roboticists and developers can determine the range of robotic applications.
Recently, a prototype, called AR2-D2, demonstrated the possibility of training robots without the need for a physical robot (Duan et al., 2023). AR2-D2 enables users to capture videos of themselves manipulating objects in various environments using smartphones or tablets. Furthermore, it has proven that the quality of data gathered through AR visualization is comparable with data collected using traditional approaches with real robots. However, AR2-D2 has only been evaluated on simple, short-horizon tasks such as pressing a computer mouse or picking up a plastic bowl. Ecologically valid user tasks, such as cleaning a table, require longer interactions between the robot and its environment. Additionally, the system’s usability has not been analyzed, which is crucial for understanding the usability and utility for every user.
We introduce EVE, an iOS app that allows everyday users to train robots with human-centric AR visualizations and features, even if they do not have access to a robot. To collect demonstrations, users can specify waypoints for the trajectory with their fingers or make the AR robot track the hand in real time. During collection, users can visually inspect the real world to avoid obstacles, modify the collected trajectory by reverting to a previous waypoint, and verify the collected trajectory by replaying the task with the AR robot.
To understand the opportunities and challenges of AR-based demonstration collection, we conducted a formative study with 10 participants with varying experience in robotics, using AR2-D2 as the baseline system and three new AR visualizations. We identified six usability challenges users faced and implemented seven additional features to improve EVE’s accuracy, flexibility, and verifiability.
We conducted a user study with 14 participants to compare EVE with state-of-the-art demonstration collection interfaces on three typical tabletop tasks. EVE achieved the highest task success rate and performed comparably to kinesthetic teaching in task completion time, system usability, motion intent communication, user enjoyment, and overall preference ($mean_{p}=0.30$). Participants customized their workflows with multiple visualizations and features to collect demonstrations for complex, long-horizon tasks.
In summary, our contributions include:
  • EVE, an iOS app that enables everyday users to train robots using intuitive AR visualizations and features
  • Formative study that identifies opportunities and challenges for improving the usability of various demonstration collection interfaces
  • User study that demonstrates how people use diverse collection interfaces
  • A discussion on designing future AR-based demonstration collection interfaces
2. Related Work
2.1. Human Demonstrations Collection for Robotics
Imitation learning enables robots to effectively automate tasks by learning from human demonstrations (Zare et al., 2023; Wang et al., 2023; Shridhar et al., 2022). A common approach for collecting demonstration data includes kinesthetic teaching, where a human physically guides the robot through a desired trajectory (Argall et al., 2009). Despite the direct haptic feedback, this mechanism can be slow and tedious (Osentoski et al., 2010). Another popular collection interface is teleoperation, where the user controls the robot remotely using specialized controllers, such as 3D space mouse (Zhu et al., 2022), VR controllers(Ebert et al., 2021; Brohan et al., 2022; Seo et al., 2023), smartphones(Mandlekar et al., 2018), and haptic devices(Shaw et al., 2023b; Toedtheide et al., 2023), which can be costly and challenging to use due to high latency and lack of intuitiveness (Akgün and Subramanian, 2011). More recent efforts also involve capturing human expert demonstrations with physical robots using specialized hardware like ALOHA(Zhao et al., 2023), GELLO(Wu et al., 2023), DexCap(Wang et al., 2024), and UMI(Chi et al., 2024). However, these studies often prioritize data collection efficiency and model training with only a few exceptions, including minimum studies (Kapadia et al., 2017; Szafir and Szafir, 2021) on examining user preference and intuitiveness in their systems.
Prior works compared kinesthetic teaching and teleoperation on simple tasks such as manipulating cups (Fischer et al., 2016; Akgün and Subramanian, 2011). The common findings are that kinesthetic teaching is preferred over teleoperation based on collection time, ease of use, and accuracy of demonstrations. In contrast to this work, our paper compared the performance of traditional collection interfaces as well as state-of-the-art AR-based collection interfaces. Furthermore, we focus on three realistic tabletop tasks with varying levels of difficulty, leading to a more comprehensive comparison between different collection interfaces.
Most closely related to our work is AR2-D2, a system for collecting demonstrations that do not require skilled demonstrators and real robots and resituating demonstration collection outside the lab setting (Duan et al., 2023). It is an iOS app that people can use to project AR robots into the real world and collect waypoint information with their hands. The user study demonstrates that AR2-D2 is intuitive to use and requires no training in contrast to four other common demonstration collection interfaces. However, AR2-D2 did not experiment with different user interface designs and only offered one way to control and visualize the robot. In addition, it only allows the collection of waypoint data, which is not useful for training the robot to do dynamic tasks such as tossing a ball. In contrast, EVE focuses on the human-centric aspect of the demonstration collection process by providing three new AR visualizations and seven new features derived from the formative study. Furthermore, EVE enables the collection of both waypoint data and trajectory data, allowing users to select suitable collection methods based on task types.
2.2. Augmented Reality and Robotics
Traditional approaches to improving human-robot interaction, such as through robots’ movements (Dragan et al., 2013; Mainprice et al., 2010), gestural motion (Huang and Mutlu, 2013; Chidambaram et al., 2012), and gaze outputs (Mutlu et al., 2006; Kalegina et al., 2018), are limited due to the challenge of presenting expressive physical feedback that goes beyond internal capabilities. In contrast, AR provides visual feedback in one’s line of sight, opening up exciting new opportunities for human-robot interaction research. A recent taxonomy of augmented reality and robotics has identified five purposes and benefits of visual augmentation (Suzuki et al., 2022). We focus on analyzing works related to motion intent communication since our goal is to improve the usability of the system by making it easier for the user to understand the goal of the AR robot.
AR interfaces improve a robot’s motion intent communication through spatial information. Prior works have experimented with different designs based on factors such as path, waypoints, velocity, arrival times, etc. (Walker et al., 2018; Zollmann et al., 2014; Rosen et al., 2017). Most relevant to our work is Rosen et al., who visualized the robotic arm’s motion intent by overlaying the expected robot poses with a mixed reality head-mounted display. They evaluated the performance of the visualization against 2D displays and no visualizations and found that the system allows a human to more quickly (-62% in time) and accurately (+16% in accuracy) tell where the robot is going to move. In addition to visualizing the robot’s motion intent, they iterated on the system to allow the robot to replan an intended trajectory with MoveIt motion planner (mov, 2023) with the same start and end points, with the final trajectory chosen by the human. This has been shown to improve task completion time, precision, accuracy, usability, and task load compared to the baseline systems. Unlike this work, our application is in the form of an iOS app, and the AR visualization is applied with a phone or a tablet. This increases the potential to enable more people to collect demonstrations for training robots compared to wearable displays. Furthermore, instead of projecting the entire pose of the robot in AR, we move the AR robot directly to the goal pose due to the lack of safety concerns. Similar to how they allow the user to replan the path, we offer a redo function that allows users to reselect the trajectory they desire.
3. EVE Prototype 1
Figure 2. Prototype 1 of EVE includes three new AR visualizations. AR kinesthetic teaching enables the robot to track the user’s hand in real-time. Path history displays trajectory along collected waypoints. Invisible robot allows users to toggle the robot body off and only view a cylinder that shows the robot’s end effector position and the gripper state (green = open, red = close).
While prior work has compared common demonstration collection interfaces and various AR visualizations, there is a lack of research investigating user preferences for AR-based demonstration collection systems. Since AR2-D2 only provides one way to control and visualize the robot’s movement, we developed the first EVE prototype with three new AR visualizations for the formative user study: AR kinesthetic teaching, path history, and invisible robot (Figure 2). Below, we describe the design and implementation of the new AR visualizations.
3.1. AR Kinesthetic Teaching (AR-KT)
Prior work suggests that kinesthetic teaching outperforms teleoperation in system usability, collection time, and task success rate (Duan et al., 2023). This is primarily due to the real-time feedback and intuitiveness when manually guiding the robot with the hand. Inspired by these findings, we implemented AR-based kinesthetic teaching (AR-KT). When AR-KT is activated, the AR robot tracks the movement of the user’s hand and adjusts the position of its end effector to match the location of the user’s hand. To further enhance the naturalness of the robot’s motion, we employed inverse kinematics to calculate the joint transformations to make the robot move in synchrony with the user’s hand.
3.2. Path History
Trajectory visualization is a popular method to improve motion intent communication, as it allows users to preview the robot’s plan before execution. Previous research has shown that visualizing proposed robot motion using a mixed reality head-mounted display led to a 16% increase in accuracy and a 62% decrease in task completion time compared to a 2D display when labeling robot arm motions as either colliding or non-colliding with objects (Rosen et al., 2017). Inspired by these findings, we incorporated a virtual line trajectory using Unity’s line renderer component. The system displayed a straight green line between the robot’s end effector coordinate and the user’s hand coordinate to minimize the occlusion of the AR robot and objects. By utilizing the Jacobian matrix to relate joint velocities to end effector velocity, applying a damped least squares approach to compute joint angle updates, and iteratively updating the joint angles, our inverse kinematics implementation enables the end effector to move in approximately straight-line motions towards the desired goal pose. Thus, linear interpolation sufficiently approximates the end effector’s actual trajectory. To support real-time control and flexibility, we included a redo button that reverts the robot to the last saved waypoint and removes the corresponding trajectory. We opted for a redo functionality instead of projecting the robot’s goal pose before movement because our entire system is AR-based, ensuring safety and efficiency when moving the robot directly to the goal pose.
3.3. Invisible Robot
The AR robot’s 1:1 scale with the real Franka Panda robot often occludes the surroundings. To address this, the invisible robot visualization renders the AR robot invisible and displays a cylinder at the end effector’s position. The cylinder’s color changes based on the gripper’s state: green when the gripper is open and red when the gripper is closed (see Figure 2). The color provides users with intuitive visual feedback about the gripper’s current state without obstructing their view of the environment.
4. formative study
Figure 3. User comments about the six demonstration collection interfaces used in the formative study.
While prior work has compared various demonstration collection interfaces, there is a lack of research on demonstration collection with AR. Thus, our formative study examines the experiences and needs of end users with AR-based collection interfaces.
4.1. Method
Participants. Since EVE aims to support everyday users, we recruited 10 participants via snowball sampling. Participants were screened via a demographic questionnaire, which asked about their demographics and backgrounds, including age, gender, area of expertise, familiarity with robotics (5-point Likert scale), familiarity with demonstration collection interfaces (5-point Likert scale), and tasks they prefer the robot to help them with. We followed best practices from the HCI Guidelines for Gender Equity and Inclusivity (Scheuerman et al., 2020) to collect gender-related information. Participants ($\mu=20.7$ years old, $\sigma=1.7$ years), denoted as F01-F10, represented seven different areas of expertise and varying levels of familiarity in robotics and demonstration collection (Table A). The participants’ familiarity with robotics ranged from unfamiliar to somewhat familiar: one person was unfamiliar, five were somewhat unfamiliar, three had neutral familiarity, and one was somewhat familiar. Regarding their familiarity with the demonstration collection, six participants were unfamiliar, and four were somewhat unfamiliar.
Procedure. The formative study took place on a university campus and lasted 50 minutes. We obtained the Institutional Review Board’s approval before the formative study, and the consent form was taken in person. All sessions were voice recorded for post hoc analysis. Because we were interested in candid reactions, we did not tell participants that we created the additional AR visualizations. Each participant was compensated with food for the study.
The formative study consisted of a two-part demonstration collection with six demonstration collection interfaces. The researchers began by demonstrating how to use each interface. Then, participants were given a few minutes to try each interface and ask any technical questions. Afterward, participants were given up to five minutes to try each interface by pushing, picking, and placing various cubes. Since we are investigating the user experience with various collection interfaces, we did not design the task to simulate a real-world scenario.
In part one, participants used kinesthetic teaching, teleoperation with an HTC Vive VR controller, and the AR2-D2 system. After the participants had tried data collection with the three state-of-the-art collection interfaces, we conducted a semi-structured interview about their experience, focusing on what features they liked and disliked about each interface. Although part one does not provide direct insight into an AR-based demonstration collection system, we included it to provide participants with a more comprehensive experience with robot data collection. In part two, participants performed the same task with AR-KT, path history, and invisible robot visualizations. Aside from discussing the same questions in part one, we brainstormed additional features for the AR-based collection interface. Subsequently, participants completed a survey for System Usability Scale (SUS) scores and a form to rank the motion intent communication, user enjoyment, and preference for the AR2-D2 system and the new AR visualizations. Finally, we explained to the participant how the participant’s data would be used to investigate the design for AR-based demonstration collection interfaces and answered any relevant questions.
Analysis. We analyzed two sources of data: interview transcripts and the post-task questionnaires. For the qualitative data, we used reflexive thematic coding to analyze the usability challenges faced by the participants with the current AR-based demonstration collection system. The first author, who facilitated all user study sessions, created an initial codebook by reviewing study transcripts. The entire team then collaboratively iterated on the codebook while checking for bias and coverage. With a final codebook consisting of 15 codes, the team discussed the resulting themes. While the formative study focuses on participants’ reactions to AR-based collection interfaces, we summarized and selected several participants’ quotes regarding what they liked and disliked about each collection interface (Figure 3). For SUS scores, we converted survey responses, which are on a scale of 0-40 when summed, to a range between 0-100(Brooke, 1996)111(((Q1 + Q3 + Q5 + Q7 + Q9) - 5) + (25 - (Q2 + Q4 + Q6 + Q8 + Q10))) * 2.5. We then conducted a Friedman test as an omnibus test with an appropriate number of Wilcoxon signed-rank tests corrected with Holm’s sequential Bonferroni procedure for statistical significance. See Table 1 for a summary of quantitative results from the post-task questionnaires.
AR2-D2 AR-KT Path History Invisible Robot
Usability (SUS)
66.3 (20.3) 67.3 (12.3) 70.0 (13.0) 68.3 (15.7)
Motion Intent (rank)
2.9 (1.2) 2.4 (1.0) 1.9 (0.9) 2.8 (1.3)
Enjoyment (rank)
3.5 (1.0)^
2.4 (1.0)
2.0 (0.7)^
2.1 (1.3)
Preference (rank)
3.5 (1.0)^
2.7 (0.8)
1.5 (0.7)^
2.3 (1.1)
Table 1. Mean and standard deviation of usability, motion intent communication, enjoyment, and preference for different AR visualizations. Usability is rated on a scale from 0 to 100, with higher scores indicating better usability. Motion intent communication, user enjoyment, and overall user preference are ranked from 1 to 4, with lower ranks indicating better performance. For statistical significance, one caret (^) is $p<0.05$.
4.2. Findings
A Friedman test found no statistically significant differences between the four AR visualizations in the SUS scores ($\chi^{2}(3)=2.6$, $p=0.5$) and rankings of motion intent communication ($\chi^{2}(3)=3.7$, $p=0.3$). The path history visualization significantly outperformed the baseline AR2-D2 system in user enjoyment ($p=0.0059$) and preference ($p=0.0059$). Notably, the path history visualization achieved the highest median scores and rankings across all measured variables: SUS ($\mu=70.0$, $\sigma=13.0$), motion intent communication ($\mu=1.9$, $\sigma=0.9$), user enjoyment ($\mu=2.0$, $\sigma=0.7$), and preference ($\mu=1.5$, $\sigma=0.7$). These findings suggest that participants found the path history visualization the most usable, enjoyable, and preferable among the four tested AR visualizations, even though it did not significantly outperform the others in all aspects. Interestingly, these results contradicted our hypotheses. Given the benefits of direct manipulation and real-time visual feedback, we had expected the AR-KT visualization to achieve the highest scores on system usability and motion intent communication compared to other visualizations. The contradiction may be due to the AR robot’s inability to accurately track the hand when it moves quickly. We also anticipated that the invisible robot visualization would receive the lowest scores on system usability and motion intent communication due to the lack of joint movement information. The contradiction could be attributed to the approximate straight-line motion of the AR robot, which allows users to easily predict the movement of the joints.
In addition, we distilled the challenges users faced with the current AR visualizations into the following six based on the semi-structured interviews:
  1. C1
    Ambiguous spatial position of the hand and the robot: Four participants (F03, F04, F08, F10) expressed difficulty interpreting the 3D environment via a 2D iPad screen. The lack of depth perception made it challenging for users to accurately gauge the relative positions of their hands and the AR robot. This ambiguity led to inaccuracies in the collected demonstrations. F03 also complained about the hand ”passing through” the robot, resulting in an unnatural collection.
  2. C2
    Lack of knowledge of joint constraints: Five participants (F04, F05, F06, F08, F10) reported a lack of information about the joint limits. This problem, coupled with the lack of depth perception via the iPad screen, causes the robot to move unexpectedly when the specified point with the hand is out of the joint limit.
  3. C3
    Imprecise trajectory visualization: The path history visualization displays a straight green line between the robot’s end effector coordinate and the user’s hand coordinate. Although three participants (F05, F06, F10) have asserted that the trajectory visualization aids their understanding of the robot’s motion intent, three other participants (F02, F04, F07) have proposed to include a more accurate trajectory.
  4. C4
    Absence of feedback regarding the collection efficacy: Five participants (F01, F04, F08, F09, F10) complained about the lack of collision detection, which prevented them from verifying whether the collected trajectory was feasible. F01 claimed that while AR is ”useful and fun,” it does not make sense for the AR robot to pass through all the obstacles.
  5. C5
    Obstructive robot design: The AR robot’s 1:1 scale correspondence with the physical Franka Panda robot frequently resulted in the occlusion of relevant objects within the scene. Four participants (F01, F05, F07, F09) reported that the AR robot’s size significantly impeded the data collection. Specifically, F01 expressed frustration regarding the robot’s obstruction of the line, while F09 noted that the robot’s presence obscured a substantial portion of the iPad’s display.
  6. C6
    Inconsistent hand tracking: Inconsistencies in hand tracking emerged as a significant concern among participants. Five individuals (F01, F02, F03, F06, F08) highlighted the presence of inaccuracies and latency in the hand tracking system employed in the AR environment. Furthermore, F02 noted that the gray line renderer, utilized in the AR2-D2 condition to visualize the tracked coordinates of the user’s fingers, proved to be visually distracting.
5. EVE Prototype 2
Figure 4. Overview of evaluation user study and prototype 2 of EVE. Left: the study involved three common tabletop tasks: toggling a switch, sorting food, and sweeping the table. These tasks were utilized to evaluate the user study with various baseline interfaces. To ensure a fair comparison of performance across these interfaces, the positions of all target objects and desired actions were fixed. Right: this section presents a sequence of frames from a sampled rollout performed by users during the evaluation. It illustrates how participants utilized the visualization and features of EVE prototype 2 to successfully complete the required demonstrations for the given task.
Informed by formative study findings and our own experiences using EVE, we created a second EVE prototype with seven additional system changes to address the usability challenges. We describe these advancements below:
5.1. Hand Position Projection
To communicate the position of the tracked hand pose in 3D space clearly (C1), a green line is generated on the surface. We have tried to attach the green line directly to the hand but found increased visual distraction and no improvement in position interpretation. To enhance motion intent communication and reduce potential collisions, the length and width of the projected line were adjusted to match the dimensions of the robot’s end effector. This improved representation of the hand position also helps prevent users from specifying coordinates that are clearly beyond the robot’s joint limits (C2).
5.2. Joint Constraints Signifier
To address user concerns about joint constraints in AR2-D2 and EVE prototype 1 (C2), we followed the joint space limits provided in the Franka Emika Panda robot specifications to define a range of reachable coordinates based on the instantiation point of the AR robot. When the user’s hand is positioned outside the reachable range, all the app interface buttons turn red (otherwise white) and are deactivated, preventing the user from changing the robot’s state.
5.3. Realistic Trajectory
Three participants (F02, F04, F07) indicated that the inaccurate, straight-line trajectory visualization did not effectively communicate the robot’s motion intent ((C3)). In response, the system was modified to display the precise, curved trajectory of the end effector. Moreover, Path History is now the default visualization for EVE, as it demonstrated superior performance compared to other AR visualizations in the formative user study.
5.4. Trajectory Replay
The absence of feedback regarding collision detection and path feasibility significantly hinders the system’s usability (C4). However, planning a trajectory for the AR robot to avoid physical obstacles often requires constructing a depth map of the world, which can be computationally expensive for the iPad. After brainstorming potential solutions with the formative study participants, we added a replay feature demonstrating the collected trajectory and changes in the gripper state using a partially transparent robot. This feature makes it easier to determine whether the robot collided with a physical object during collection. This implementation was also inspired by some participants’ preferences (F02, F03) for the replay feature available in kinesthetic teaching. To integrate trajectory visualization and enable replay for AR-KT, we experimented with various numbers of waypoints and decided to use 20 waypoints for the trajectory collected with AR-KT.
5.5. Dynamic Camera View
To address the issue of the AR robot occupying a significant portion of the screen (C5), a formative study participant (F05) proposed enabling the user to move the camera to inspect the 3D world. However, collecting demonstrations from multiple camera angles is not ideal because it is crucial to ensure that the hand-eye calibration between the robot action and the 3D voxels is aligned for a successful transfer to real-world deployment. To address this, we save the initial camera extrinsic values after the robot has been instantiated and make the buttons disappear when the current extrinsic values differ from the saved values by more than a small threshold. This approach allows the user to freely inspect the scene from multiple angles while ensuring the demonstration is collected from a consistent camera angle. This feature also addresses the lack of information about joint constraints (C2) by allowing users to visually inspect the robot’s state and avoid specifying potentially unreachable coordinates.
5.6. Accurate Gripper Control
In AR2-D2, a gray line is rendered between the thumb and index fingers to enable users to toggle the gripper opening by adjusting the distance between the fingers. Due to limitations in the iPad camera and ARKit’s HandTrackingProvider, hand tracking is inconsistent, and toggling the gripper state is extremely unpredictable (C6), as reported by five participants from the formative study (F01, F02, F03, F06, F08). To maintain the system’s ubiquity, we did not modify the hardware. Instead, we addressed the issue of the flashing gray line renderer by removing it and adding a button to toggle the gripper opening accurately. This change ensures a more reliable and predictable user experience when controlling the gripper state.
5.7. Robot Instantiation Reselection
Our experience using the system and the feedback from one participant (F09) suggest that the inability to change the instantiation point of the robot interferes with the system’s usability, as we need to quit and restart the app every time we instantiate the robot at an undesired place. We changed the robot instantiation process into two steps to address this issue. First, a green cylinder appears on the screen, signifying where the robot will spawn. The user is then prompted with two choices: 1) instantiate the robot at the position indicated by the cylinder, or 2) reselect a position for the green cylinder. The ability to reselect the instantiation point for the AR robot greatly increases the system’s usability.
6. evaluation user study
We conducted a user study to evaluate the effectiveness of EVE compared to baseline interfaces for three common tabletop tasks. These tasks were designed based on the results of an extensive survey from BEHAVIOR-1K (Li et al., 2024) asking participants, ”What do you want robots to do for you?” We included obstacles in all tasks to simulate real-world environments where tables are often cluttered. The three tasks were as follows:
  • Toggle switch: Participants guided the robot to push a switch and turn on a light bulb.
  • Sort food: Participants were required to grab a melon and place it inside a yellow bowl, and then grab a steak and put it into a basket. This task simulated the sorting of different items into specific locations.
  • Sweep table: Participants used a sponge to sweep beans to a location marked by tape. This task simulated the cleaning of debris from a table.
The setups for these tasks are illustrated in Figure 4.
6.1. Method
Participants. To ensure a diverse area of expertise, we recruited 14 participants via snowball sampling. Participants were screened using the same demographic questionnaire from the formative study. We followed best practices from the HCI Guidelines for Gender Equity and Inclusivity to collect gender-related information. Participants ($\mu=22.6$ years old, $\sigma=2.6$ years), denoted as E01-E14, represented 11 different areas of expertise and varying levels of familiarity in robotics and demonstration collection (Table A). The participants’ familiarity with robotics ranged from unfamiliar to familiar: three people were unfamiliar, six people were somewhat unfamiliar, three people had neutral familiarity, and two people were familiar. Regarding their familiarity with the demonstration collection, eight people were unfamiliar, three people were somewhat unfamiliar, two people had neutral familiarity, and one person was somewhat familiar. Each participant performed one to three tasks, resulting in a total of 30 collections (10 collections per task). None of the participants participated in the formative study.
Procedure. The in-person laboratory study took place on a university campus, and each task lasted 40 minutes. The consent form was taken in person. Because we were interested in candid reactions, we did not tell participants that we created EVE. Each participant was compensated with food for the laboratory study.
We began by providing a short tutorial for all the collection interfaces (i.e., kinesthetic teaching, teleoperation with an HTC Vive VR controller, AR2-D2, and EVE) and informed them of the goal of the task(s). We did not demonstrate how to accomplish the task(s) because we did not want to influence the trajectory collected by the participants. After a 30-second practice session before each collection, participants were asked to complete the task in three minutes. We recorded each collection’s task success rate and remaining time (three minutes - collection time). We allowed the participant to retry the collection if a collision occurred and the collection time had not reached three minutes. After the participants had completed the collection with all the collection interfaces, we asked them to complete the SUS survey and a form to rank the motion intent communication, user enjoyment, and overall preference for all collection interfaces.
Analysis. We analyzed the results from the SUS survey and ranking forms with the same method used for the formative study result analysis. See Table 2 for a summary of quantitative results from the post-task questionnaires. We calculated the overall task success rate and the mean and standard deviation of the remaining time for each collection interface.
6.2. Findings
EVE achieved the highest task success rate and performed comparably to kinesthetic teaching in task completion time, system usability, motion intent communication, user enjoyment, and overall user preference ($mean_{p}=0.30$). Kinesthetic teaching attained the highest SUS score and ranked the highest in motion intent, user enjoyment, and overall preference. However, there was no statistically significant difference between EVE and kinesthetic teaching, suggesting that EVE may be a viable substitute for collecting trajectory data. This might be because of the ability to clearly inspect the real world before moving the robot, a feature not present in AR2-D2 and teleoperation. Our results confirm the findings of previous work (Wrede et al., 2013; Mandlekar et al., 2018), which showed that kinesthetic teaching outperforms VR controllers in all measurements. EVE outperformed AR2-D2 in both demonstration collection success rate and remaining time. In all tasks performed by the participants, EVE surpassed AR2-D2 in SUS score, motion intent, enjoyment, and overall preference ranking.
Figure 5. The mean success rate (%) for the evaluation user study, which encompassed three tasks with 10 trials each ($D=10$), indicates that the EVE achieved a 100% success rate across all tasks. This performance surpasses that of the second-highest interface, kinesthetic teaching, by an average margin of 10% across the tasks.
Figure 6. The average remaining time in seconds and its standard deviation taken to successfully collected one demonstration for each tasks. EVE’s performance in terms of time efficiency is comparable to that of kinesthetic teaching, with a narrow average difference of 5.1 seconds across three tasks.
Overall Kinesthetic Teaching Teleoperation AR2-D2 EVE
Usability (SUS)
86.0 (17.0)
27.3 (17.8)*
59.0 (20.5)*
77.8 (17.0)
Motion Intent (rank)
1.4 (0.7)
3.3 (0.8)*
3.4 (0.7)*
1.9 (0.8)
Enjoyment (rank)
1.5 (0.8)
3.2 (1.0)*
3.3 (0.7)*
2.0 (0.8)
Preference (rank)
1.7 (0.9)
3.4 (0.9)*
3.1 (0.7)*
1.8 (0.7)
Toggle Switch Kinesthetic Teaching Teleoperation AR2-D2 EVE
Usability (SUS)
83.8 (18.0)
31.3 (21.5)*
66.0 (20.2)*
82.8 (13.9)
Motion Intent (rank)
1.4 (0.7) 3.4 (1.0)
3.2 (0.6)*
2.0 (0.8)
Enjoyment (rank)
2.0 (1.1) 3.0 (1.2)
3.1 (0.9)*
1.9 (1.0)
Preference (rank)
2.2 (1.0) 3.3 (1.1)
2.9 (0.9)*
1.6 (0.8)
Sort Food Kinesthetic Teaching Teleoperation AR2-D2 EVE
Usability (SUS)
88.0 (17.1)*
26.5 (15.5)*
54.0 (20.6)*
71.0 (17.5)
Motion Intent (rank)
1.3 (0.5) 3.1 (0.7)
3.7 (0.5)*
1.9 (0.7)
Enjoyment (rank)
1.2 (0.4)*
3.1 (1.1)
3.5 (0.5)*
2.2 (0.6)
Preference (rank)
1.2 (0.4)
3.5 (0.7)*
3.4 (0.5)*
1.9 (0.6)
Sweep Table Kinesthetic Teaching Teleoperation AR2-D2 EVE
Usability (SUS)
86.3 (17.5)
24.3 (17.1)*
56.8 (20.9)*
80.0 (18.7)
Motion Intent (rank)
1.6 (0.8) 3.3 (0.8)
3.3 (0.8)*
1.8 (0.8)
Enjoyment (rank)
1.4 (0.7)
3.5 (0.8)*
3.2 (0.6)*
1.9 (0.7)
Preference (rank)
1.6 (0.8)
3.5 (0.8)*
3.1 (0.7)*
1.8 (0.8)
Table 2. Mean and standard deviation of usability, motion intent communication, enjoyment, and overall preference for different demonstration collection interfaces. Usability is rated on a scale from 0 to 100, with higher scores indicating better usability. Motion intent communication, enjoyment, and overall preference are ranked from 1 to 4, with lower ranks indicating better performance. One asterisk (*) indicates statistical significance ($p<0.05$) when compared to the EVE system.
Below are some notable comments from users based on their observations of different collection interfaces used in three tasks during the evaluation study:
Toggle Switch. This task involves operating an augmented robot gripper to navigate around potential obstacles and accurately activate a switch. In kinesthetic teaching, users appreciated the method’s simplicity and precise effector control, facilitating movement to the desired location. However, this system does not support the modification of previously established waypoints. For teleoperation and AR2-D2, many users reported failures, attributing them to collisions with scene distractors and the systems’ inability to provide a comprehensive scene view due to robotic arm obstructions. Specifically, teleoperation users struggled with accurately positioning the robot end-effector to engage the switch, hindered by a lack of haptic feedback. Conversely, in our EVE system, users found the AR-KT features significantly improved control during the final positioning of the gripper, and the invisible robot mode provided a clear view to assess potential collisions with distractors.
Sort Food. This task required users to control a robot to perform a series of long-horizon actions, including picking up and placing multiple objects. In kinesthetic teaching, users found it straightforward to maneuver the robot for pick-and-place actions, praising its intuitiveness and ease of use. However, in teleoperation, the robot’s obstruction of the end-effector’s view complicated precise control, leading to difficulty in achieving the ideal grasp pose and frequent minor errors, often resulting in objects being in ungraspable states. In AR2-D2, users struggled due to inadequate visualization of potential collisions between the robot arm and table distractors. Conversely, the EVE system was well-received, with users appreciating the path history for tracking long-horizon actions and the redo feature for correcting waypoints. The AR-KT functionality was highlighted for its precise control, allowing for accurate object placement and enhanced gripper manipulation.
Sweep Table. In this task, users needed precise control to grasp onto the sponge and execute a series of sweeping actions. Kinesthetic teaching was favored for its simplicity, allowing users to manipulate the robot’s end-effector directly. For the teleoperation system, users encountered difficulties due to latency and the precision required to control the end-effector with a VR controller. They reported struggling to maintain a consistent direction of movement when controlling the end-effector, often leading to collisions between the end-effector and the table. With AR2-D2, the challenge was also in using fixed forward control to linearly move the end-effector along one axis, compounded by unreliable tracking of the desired gripper state. Conversely, the EVE system provided enhanced precision through its AR-KT features, offering accurate gripper control that ensured the gripper maintained the state desired by the user.
7. discussion
In this section, we reflect on our findings from the development and evaluation of EVE. We also discuss future opportunities for research exploring AR-based demonstration collection interfaces for robotics.
7.1. Implications
Personalized Robot Training in Everyday Environments. By allowing users to easily capture videos and manipulate objects in various locations outside of traditional laboratory settings, EVE enables individuals to expose robots to a wide range of objects and environments that are relevant to their specific needs and preferences. As the cost of robot hardware decreases and household robots become more prevalent, EVE puts the power of robot training directly into the hands of users, allowing them to collect personalized demonstrations for specific household tasks and fine-tune their robots to perform these tasks effectively in their unique home environments.
Integrating with Other Types of Robots. EVE allows users to quickly adapt the system to work with different robot platforms. Because each component of the program is separated, by replacing the robot prefab in Unity with a different robot model, such as the UR5e, users can collect demonstrations with a new AR robot. Furthermore, since the waypoint data collected by EVE is a standard input to multiple motion planners, EVE enables the training of manipulation policies for different robots without modifying the collected data. This flexibility empowers individuals to extend the applicability of EVE to a wide range of robotic systems, making robot training accessible to a broader audience.
7.2. Limitations
Multi-View Data Collection. EVE is limited to single-viewpoint data collection to maintain accurate hand-eye calibration between the robot’s actions and the corresponding 3D voxel representation, which is essential for successful real-world deployment. However, multi-view data collection could be helpful in scenarios where objects are occluded from one perspective but visible from another, or when the demonstrator needs to maintain a more comfortable posture by collecting trajectories from different angles. Future work could explore the use of simultaneous localization and mapping (SLAM) techniques (Durrant-Whyte and Bailey, 2006) to track the iPad LiDAR’s pose and build a 3D map of the workspace, allowing the demonstrator to collect trajectories relative to this 3D map.
Bimanual Robotic Manipulation. Many real-world tasks, such as washing dishes, ironing clothes, and assembling furniture, require bimanual manipulation. However, collecting demonstrations for bimanual tasks using EVE is challenging due to the inability to press buttons during collection. Future research could explore the use of mixed reality head-mounted displays for demonstration collection. With the increasing availability of novel head-mounted displays, collections using these devices might become more affordable and ubiquitous. Alternative input modalities, such as gaze and speech, could be used to trigger different functions instead of clicking buttons.
Collection for Mobile Robot. The real world also involves tasks that require mobile robots, such as cleaning the floor and watering plants. Extending EVE to support demonstration collection for mobile robots presents an exciting direction for future work. To achieve this, researchers can integrate a 3D model representing the mobile robot into Unity and develop functionalities to control the robot’s movement through intuitive touch gestures. The system can then leverage the device’s camera and inertial measurement unit (IMU) sensors to construct a detailed map of the environment while simultaneously tracking the robot’s pose, enabling data collection for mobile robots.
Hardware. The current system’s hand-tracking capabilities are not fully optimized due to hardware constraints. The LiDAR technology employed in the iPad occasionally results in inaccuracies in the end effector’s movement and insufficient anchoring of the robot base on the surface. To address this issue and improve tracking accuracy and robustness, future work can implement sensor fusion techniques by combining data from multiple sensors, such as the iPad’s camera, inertial measurement unit (IMU), and LiDAR scanner.
8. conclusion
We introduced EVE, a system that democratizes robot data collection with human-centric AR visualizations. Informed by our formative study ($N=10$), EVE enables end users to collect, change, and verify valuable trajectory data to train robot manipulation for personalized tasks. Our evaluation user study ($N=14,D=30$) demonstrated improved task success rate, completion time, system usability, motion intent communication, user enjoyment, and user preference against state-of-the-art collection interfaces. We hope EVE will catalyze future work in AR-based demonstration collection for robotics and expedite the incorporation of robots into daily lives.
Appendix A Study Participants Demographics
# f e Area of Expertise Familiarity (Robotics) Familiarity (Demonstration Collection) Tasks
1 F01 -
Marketing
Somewhat unfamiliar
Somewhat unfamiliar
Chores, remind about deadlines
2 F02 -
Math
Neutral
Somewhat unfamiliar
Chores
3 F03 -
Food Systems
Somewhat unfamiliar
Unfamiliar
Cooking
4 F04 -
Applied Math
Unfamiliar
Unfamiliar
Cooking, driving
5 F05 -
Math
Somewhat unfamiliar
Unfamiliar
Manage files
6 F06 -
Music
Neutral
Unfamiliar
Grab packages, sort folders
7 F07 -
Applied Math
Somewhat familiar
Unfamiliar
Fetch items
8 F08 -
Computer Science
Somewhat unfamiliar
Somewhat unfamiliar
Cooking, driving, grab packages
9 F09 -
Statistics
Somewhat unfamiliar
Unfamiliar
Chores, teaching
10 F10 -
Computer Science
Neutral
Somewhat unfamiliar
Assemble car components
11 - E01
Music
Unfamiliar
Unfamiliar
Grab packages
12 - E02
Electrical Engineering
Somewhat unfamiliar
Unfamiliar
Chores
13 - E03
Statistics
Unfamiliar
Unfamiliar
Fetch items
14 - E04
Business
Somewhat unfamiliar
Somewhat unfamiliar
Chores
15 - E05
Computer Science
Somewhat unfamiliar
Unfamiliar
Fold clothes
16 - E06
Electrical Engineering
Neutral
Somewhat unfamiliar
Cooking
17 - E07
Education
Somewhat unfamiliar
Unfamiliar
Teaching
18 - E08
Informatics
Somewhat unfamiliar
Unfamiliar
Grab packages
19 - E09
Genetics
Neutral
Unfamiliar
Cleaning
20 - E10
Computer Science
Neutral
Neutral
Sort folders
21 - E11
Informatics
Somewhat unfamiliar
Somewhat unfamiliar
Grab packages
22 - E12
Math
Unfamiliar
Unfamiliar
Teaching
23 - E13
Robotics
Familiar
Neutral
Fold clothes
24 - E14
Mechanical Engineering
Familiar
Somewhat familiar
Factory inspection, fetch items
Table 3. Participant demographics in formative and evaluation study. In total, there were 24 participants overall (#). f denotes the participant IDs for the formative study, e denotes the participant IDs for the evaluation study.